Week Of August 31 September 14 September 2 Pints Used 350 370 410 378 366 374 October 5 a) The forecasted demand for the week of October 12 using a 3-week moving averagepints (round your response to two decimal places). b) Using a 3-week weighted moving average, with weights of 0.10. 0.35, and 0.55, using 0.55 for the most recent week, the forecasted demand for the week of October 12pints (round your response to two decimal places and remember to use the weights in appropriate order the largest weight applies to most recent perod and smalest weight applies to aldesf peniod.) c) f the forecasted demand for the week o August 31 is 350 and α-0.25, using exponential smoothing, develop the forecast or each of the weeks with the known demand and the forecast for he week of October 12 round your responses to two decimal places). Week Of Pints Used Forecast for this Date 350 350.00 365.00 August 31 350 370 410 378 366 374 September 28 October5 October 12 Enter your answer in each of the answer boxes.
Answer #1

Ans a)

A 3-week moving average forecast uses the average value of actual sales/demand of the immediately previous three periods.

The forecast for October 12 = (378 + 366 + 374) / 3

= 372.67 pints

Ans b)

A 3-week weighted moving average forecast uses the weighted average value of actual sales/demand of the immediately previous three periods.

Weight for October 5 = 0.55

Weight for September 28 = 0.35

Weight for September 21 = 0.10

The weighted forecast for October 12 = (374 * 0.55 + 366 * 0.35 + 378 * 0.10)

= 371.6 pints

Ans c)

To find the exponential smoothened forecast, we can do the calculations in Excel.

Forecast for 31 August = 350 pints

\alpha = 0.25

Formulas used

From the above screenshot of the spreadsheet, the exponential forecast -

for Sep 21 = 368.75 pints

for Sep 28 = 371.06 pints

for Oct 5 = 369.80 pints

for Oct 12 = 370.85 pints